2a. Calculate ΔG∘rxn at 25∘C for a redox reaction with n = 4 that has an equilibrium constant of K = 5.1×10−2.

2b. Calculate E∘cell at 25∘C for a redox reaction with n = 4 that has an equilibrium constant of K = 5.1×10−2.
Answer #1

a)

T= 25.0 oC

= (25.0+273) K

= 298 K

use:

ΔGo = -R*T*ln Kc

= - 8.314*298* ln(5.1*10^-2)

= 7373.08 J/mol

= 7.37 KJ/mol

Answer: 7.37 KJ/mol

b)

use below formula for finding Eo from K

Eo = (2.303*R*T)/(n*F) log Kc

At 25 oC or 298 K, 2.303*R*T/F = 0.0592

So, Eo = (0.0592/n)*log Kc

Eo = (0.0592/4)*log(5.1*10^-2)

Eo = -1.913*10^-2 V

Answer: -1.91*10^-2 V